a 150 gram ball at the end of a string is swinging in a horizontal circle of radius 1.15. the ball takes 5 seconds to complete 20 revolutions. a. what is the time for 1 revolution of ball? b. what is the tangential speed of the ball? c. what is the tension in the string?
Answer #1

Here,

a) time period = time taken/number of revs

time period = 5/20

time period = 0.25 s

the time period is 0.25 s

b) tangential speed of the ball = 2pi * r/time period

tangential speed of the ball = 2pi * 1.15/(0.25)

tangential speed of the ball = 28.9 m/s

tangential speed of the ball is 28.9 m/s

c)

tension in the string = m * v^2/L

tension in the string = 0.150 * 28.9^2/1.15

tension in the string = 108.9 N

the tension in the string is 108.9 N